Marshall Company is issuing four-year bonds with a coupon rate of 4.5 percent and semiannual coupon payments. If the current market rate for similar bonds is 9.2 percent, what will be the bond price? Round to 2 decimal places.
Answer #1

Price of a bond is mathematically represented as: 1 1-- (1 + i) PEC. - +- м. (1 + i) where P is price of bond, with periodi

M = $1000, n = 4 * 2 = 8 semi-annual periods, C = 4.5% * $1000/2 = $22.50 (semi-annually), i = 9.2%/2 = 4.6% (semi-annual)

1000 P=22.50* 1-71-0.0465 2016 -(1+0.046)

P = $147.80 + $697.82

P = $845.63